OpenStudy (anonymous):

A bulldozer moves 640 lb of dirt in 4/5h. Complete the unit rate.

OpenStudy (anonymous):

what r the answer choices

OpenStudy (anonymous):

there are none

OpenStudy (yuliacosta111):

640 (ib) --------->(4/5) hour X ----------> (1) hour x=(1)*(640)/(4/5)=640/4*5=160*5= 800